HUBBLE SPACE TELESCOPE DAILY REPORT #2535 PERIOD COVERED: 0000Z (UTC) 01/07/00 - 0000Z (UTC) 01/10/00 Daily Status Report as of 010/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of S/C 8518 (Vehicle Disturbance Test) The Vehicle Disturbance Test {VDT} is used to characterize the HST dynamic response due to environmental disturbances during extended gyro hold intervals. Disturbances are caused by on-board mechanism motion and transient effects resulting from thermal strain. These disturbances excite the natural modes of HST structural appendages and produce vehicle dynamic response as measured by the rate gyros. Disturbance signatures are compared with signatures taken after previous servicing missions to identify changes in mechanisms or structures since the last VDT. The disturbances observed will be used as the nominal on-orbit disturbances in pointing control simulations until the next VDT is run. At least 5 orbits at +V3 sunpoint {with 90-degree solar array angle} and 3 orbits at -V1 sunpoint {with 0-degree solar array angle} are required for this test. Using two different solar array orientations allows additional identification of disturbances induced by the solar arrays. Low bandwidth attitude control gains are used to configure the HST for "nearly open-loop" dynamic response during the test. This configuration allows observation of HST disturbance response that is nearly independent of the affects of the pointing control law. No science instruments are used during this test. The test completed with no reported problems. 1.2 Completed Five Sets of FGS 8532 (White Dwarf and M-Stars in the Solar Neighborhood; Revealing Binarity) FGS-1 was used to conduct a high angular resolution survey with FGS1r to detect binarity among nearby M dwarfs and white dwarfs. The observations completed with no reported problems. 1.3 Completed STIS/CCD 8507 (CCD Darks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain weekly updates to the hot pixel list. The proposal completed as planned. 1.4 Completed Two Sets of STIS/CCD 8508 (CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to obtain updates to the weekly superbiases. The proposal was completed as planned. 1.5 Completed Four Sets of STIS/MA2 8509 (NUV MAMA Darks) The Space Telescope Imaging Spectrograph (MA2) was used to determine the health of the FUV MAMA detector by monitoring the background count rate. The observations completed with no reported problems. 1.6 Completed Four Sets of WF/PC-2 8492 (UV Monitor for -88C Cooldown) The WF/PC-2 was used to conduct the UV monitor after the first cooldown to -88C. Throughput measurements are also made using UV filters and UV crossed with long-pass filters to determine red leak contribution. -These are used to indicate any drop in throughput at Lyman alpha, which could be due to contamination in the pick-off mirror. The proposal completed nominally. 1.7 Completed S/C 8516 (Gyro/FHST Alignment) The purpose of this proposal is to calibrate the scale factors and alignments of the gyros that were replaced during SM3A. There were no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 13 Successful: 13 Scheduled Re-acquisitions: 0 Successful: 0 2.2 FHST Updates: Scheduled: 45 Successful: 53 According to HSTAR 7469, the full maneuver update using FHSTs 1 and 2 at 009/144658Z failed, as did the following full maneuver update at 009/144943Z. 2.3 Operations Notes: Servicing Mission Observatory Verification (SMOV) activities dominated the observations conducted over this weekend. On several occasions, the SSR EDAC error counter was cleared via ROP SR-1A. Throughout the weekend, six high/low gyro bias update loads were uplinked, and five attitude reference updates were performed. A TTR was written for a period of no GCMR capability, beginning at 007/2223Z. An RGA/FHST calibration exercise was conducted from 008/0315Z to 008/1315Z (see 1.7). Resulting RGA/FHST calibration tables were uplinked at 009/1510Z. The Vehicle Disturbance Test (see 1.1) was installed and conducted successfully from 008/1835Z until 009/1210Z. During the Test, a VAP off-position type T2 slew was performed at 009/0515Z. Also during the test a realtime map and a GENSLEW were performed. The SM3A secondary mirror move was performed at 009/1750Z. The total requested move, from -2.9 microns to +1.0 microns, is 3.9 microns defocus but, since actuator step commanding has an approximate 0.60 micron resolution, an actual 4.2 micron move was accomplished. The STIS MCE-1 reset at 009/194646Z while the low voltage was on and while in an SAA interval. The STIS flight software error counter was reset at 009/2014Z per ROP NS-12. MAMA-1 will be recovered at 010/1331Z, the time of the next high voltage on commanding via the SMS. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of SMOV and normal science observations and calibrations. /CAW